# Quantum Vacuum energy as the origin of Gravity

### v2

@@ IV Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
with ${\cal G}(\mu)$ defined in (43). For the moment, let us not tie the energy scale to $\mu$ to $z(t)$ as we did in (42). It is conventional to express (50) as follows
@@
where by definition $a(t_{0})=1$ at the present time $t_{0}$, $H(t_{0})=H_{0}$, and $\Omega_{\Lambda}$ is the $\rho_{\rm vac}$} contribution. The zero curvature ($k=0)$ motivated in Section II implies $\Omega_{\rm rad}+\Omega_{\rm m}+\Omega_{\Lambda}=1$. In the $\Lambda{\rm CDM}$ model one ignores radiation, i.e. $\Omega_{\rm rad}\approx 0$, as this is a well-justified approximation during this epoch.
There is a deepening discrepancy in the $\Lambda{\rm CDM}$ model based on relatively recent astrophysical measurements referred to as the “Hubble tension”. The ideas in this article offer a potential resolution which we now present. As usual, let $H_{0}=\dot{a}/a$ be the Hubble constant today. It should be a fixed constant regardless of the manner in which it is measured. The discrepancy arises from two very different kinds of measurements. The first comes from “local” measurements based on Type Ia supernovae, where typical redshifts are relatively low, in the range z=0.020.15z=0.02–0.15 (SHOES SHOES .). The other determination of $H_{0}$ is based on Cosmic Microwave Background data at much higher redshift $z=1100$. (Planck Planck ). The two contradictory values reported are
@@
which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef . In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was proposed.
@@
$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100,$ (53)
@@
Let now turn to explicitly solving (51) in “real time”, namely with ${\cal G}(\mu(t))$ given in (42), as we did for pure vacuum energy with the result (45). Before even integrating the equation (50), one can immediately see that there are no real solutions unless $1+\widehat{b}\log a>0$. Thus there is still a minimal scale factor $a(t)>a_{\rm min}$ which is the same as in (46) and attributed to the Landau pole. This is a robust conclusion, regardless of whether $\rho_{\rm rad}$ is incorporated as the dominant form of energy in the very early Universe. The above value for $\widehat{b}$, leads to a very high value for the maximal red shift: [^17] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$. For comparison, for QED $b=4N_{f}/3$ for $N_{f}$ species of Dirac fermions, and if $\mathfrak{g}_{0}=1/137$ then $\widehat{b}=.002N_{f}$ which is about 10 times smaller than in (54) and $z_{\rm max}\approx 10^{500}$ is far beyond the Planck scale.
$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}.$ (54)
@@
To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ as a good approximation,. In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (50), let us neglect the $\Omega_{\Lambda}$ term as an approximation. Then one wishes to integrate the equation
@@
where $a(t_{0})=1$ is an initial condition. There exists a time $t_{\rm min}$ where aa=0: [^18] In the cosmology literature the convention is to shift $t\to t+t_{\rm min}$ such that $a=0$ at the shifted time $t=0$. The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assume the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
@@
where again we have imposed $a(t_{0})=1$. The argument of the $\rm erfi$} function on the LHS must be real otherwise the $\rm erfi$} function is imaginary. Thus due to the square-root branch cut $\sqrt{1+\widehat{b}\log a}$ inside $\rm erfi$}, there is a minimal value of $a(t)$ again given by (46). This minimal value of $a(t)$ occurs at a time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$, which is determined by when the LHS of the above equation is zero since ${\rm erfi}(0)=0$:
@@
Based on the above analysis of our scenario, let us comment on the possible implications for the very early Universe. Henceforth, by “early Universe” we mean around the time $t_{\rm min}$, where the expansion phase begins. Recall in our model, time $t$ actually extends to $t=$-$\infty$, and this is the earliest Universe. The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, described in the last paragraph of the last section. Namely at $t=$-$\infty$, $a(t)=\infty$ and starts to decrease until it reaches $a_{\rm min}$ at time $t_{\rm min}$, then begins to increase back to $\infty$ as $t\to+\infty$. At no time in this evolution is $a=0$, again avoiding geometric curvature singularities. There is no singular Big Bang at any time, rather at time $t_{\rm min}$ the universe is hottest since the redshift $z$ is at its maximum. The quantity $t_{0}-t_{\rm min}$ does not represent the age of the Universe, but rather the time elapsed since the scale factor was at its minimum. The age of the Universe is actually infinite, which we find appealing, since the Universe does not originate from an incomprehensible singularity at $a=0$. As described above for the case of only Vacuum energy, the time evolution of the Universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing.
@@
The details of this evolution are very sensitive to the value of $\widehat{b}$ due to the exponentials in (54). For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred from the Hubble tension. First, one sees from (60) that for small $\widehat{b}$ this amounts to a small corrections to $t_{0}-t_{\rm min}$ which is normally associated to the age of the Universe. If our estimate of $\widehat{b}$ in (54) is approximately correct, then this value for $z_{\rm max}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$. At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (42) is not necessarily valid, however if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude. This indicates that the Planck scale plays no role in determining $z_{\rm max}$, which is to say quantum gravity does not play a role.
@@
For our perspective on gravity, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model. In the standard cosmology this earlier era is thought to be described by inflation. Inflation is postulated as a way to get from the singularity at $a=0$ to a hot Big Bang, however it is not yet universally accepted as being a complete theory. Very large numbers are naturally involved since the singularity at $a=0$ formally corresponds to $z=\infty$. Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a\approx 0$, then transition to a re-heating scale to bring up the temperature to that of a hot Big Bang, and this rapid expansion makes the Universe flat and homogeneous. The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is, and the duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$, where the upper limit can even be significantly larger than $z_{\rm Planck}$. [^19] During hypothetical inflation, the redshift $z$ is not given by the formula (42), but based on the definition of the scale factor $a=1/(1+z)$. It is noteworthy that at the lower end of this range based on lower reheating temperatures, namely where inflation ends at $z_{\rm inflation}=10^{22}$, is close to our $z_{\rm max}=5\times 10^{21}$. Since this inflationary epoch does not actually exist in our model, one should question whether inflation is necessary in the standard model of cosmology. At $t=t_{\rm min}$ the Universe is already expanded since $a(t)>a_{\rm min}$, and the Universe is already flat with $k=0$. In Section II we explained why the spatial curvature $k$ is naturally zero since we view Gravity as arising from vacuum energy in flat Minkowski space. The value of $a_{\rm min}=e^{-1/\widehat{b}}$ perhaps plays a role equivalent to hypothetical inflation models since it represents roughly 50 e-foldings since $e^{1/\widehat{b}}\approx e^{50}$ for our estimate of $\widehat{b}\approx 0.02$.
V Closing Remarks

### v3

@@ VI Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
with ${\cal G}(\mu)$ defined in (55). It is conventional to express (62) as follows
@@
where by definition $a(t_{0})=1$ at the present time $t_{0}$, $H(t_{0})=H_{0}$, and $\Omega_{\Lambda}$ is the $\rho_{\rm vac}$} contribution. As for the dark-universe, the additional time dependence due to the factor $(1+\widehat{b}\log a(t))$ does not follow directly from the Einstein equations, but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (54). The Bianchi identities are only valid locally where $\mu$ is a fixed and thus Newton’s constant ${\cal G}(\mu)$ is time independent. The zero curvature ($k=0)$ motivated in Section II implies $\Omega_{\rm rad}+\Omega_{\rm m}+\Omega_{\Lambda}=1$. The above prescription (63) for dealing with the issue raised preserves the minimal scale factor $a_{\rm min}$ as we will show below. It is also a key aspect of our proposal for dealing with the Hubble tension.
In the $\Lambda{\rm CDM}$ model one ignores radiation, i.e. $\Omega_{\rm rad}\approx 0$, as this is a well-justified approximation during this epoch. There is a deepening discrepancy in the $\Lambda{\rm CDM}$ model based on relatively recent astrophysical measurements referred to as the “Hubble tension”. The ideas in this article offer a potential resolution which we now present. As usual, let $H_{0}=\dot{a}/a$ be the Hubble constant today. It should be a fixed constant regardless of the manner in which it is measured. The discrepancy arises from two very different kinds of measurements. The first comes from “local” measurements based on Type Ia supernovae, where typical redshifts are relatively low, in the range z=0.020.15z=0.02–0.15 (SHOES SHOES .). The other determination of $H_{0}$ is based on CMB data at much higher redshift $z=1100$. (Planck Planck ). The two contradictory values currently reported are
@@
which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef ; Riess2 . In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was advocated.
@@
$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100~~\Longrightarrow~~\widehat{b}\approx 0.02$ (65)
@@
Let now turn to explicitly solving (63) in “real time”, namely with ${\cal G}(\mu(t))$ given in (54), as we did for pure vacuum energy with the result (57). Before even integrating the equation (62), one can immediately see that there are no real solutions unless $1+\widehat{b}\log a>0$. Thus there is still a minimal scale factor $a(t)>a_{\rm min}$ which is the same as in (58) and attributed to the Landau pole. This is a robust conclusion, regardless of whether $\rho_{\rm rad}$ is incorporated as the dominant form of energy in the very early Universe. The above value for $\widehat{b}$, leads to a very high value for the maximal red shift and temperature: [^19] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$. For comparison, for QED $b=4N_{f}/3$ for $N_{f}$ species of Dirac fermions, and if $\mathfrak{g}_{0}=1/137$ then $\widehat{b}=.002N_{f}$ which is about 10 times smaller than in (66) and $z_{\rm max}\approx 10^{500}$ is far beyond the Planck scale.
$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}~~~\Longrightarrow~~~T_{\rm max}\approx 10^{22}\,K.$ (66)
@@
To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ is a good approximation. The scale factor $a(t)$ can be solved numerically starting from (63). In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (62), let us neglect the $\Omega_{\Lambda}$ term as an approximation, since the resulting equation can be solved analytically and leads to more transparent conclusions. Then one wishes to integrate the equation
@@
where $a(t_{0})=1$ is an initial condition. There exists a time $t_{\rm min}$ where aa=0: [^20] In the cosmology literature the convention is to shift $t\to t+t_{\rm min}$ such that $a=0$ at the shifted time $t=0$. The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assumes the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
@@
where again we have imposed $a(t_{0})=1$. The argument of the $\rm erfi$} function on the LHS must be real otherwise the $\rm erfi$} function is imaginary. Thus due to the square-root branch cut $\sqrt{1+\widehat{b}\log a}$ inside $\rm erfi$}, there is a minimal value of $a(t)$ again given by (58). It is not difficult to check that the same conclusion is reached in a radiation dominated era where the $1/a^{3}$ term is replaced by $\Omega_{\rm rad}/a^{4}$. This minimal value of $a(t)$ occurs at a time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$, which is determined by when the LHS of the above equation is zero since ${\rm erfi}(0)=0$:
@@
Based on the above analysis, let us comment on the possible implications for the very early universe in our model. Henceforth, by “early universe” we mean around the time $t_{\rm min}$, where the universe is hottest and the expansion phase begins. Recall in our model, time $t$ actually extends to $t=$-$\infty$, and this is the earliest universe. The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, i.e. the dark-universe  described in the last paragraph of the last section, and this justifies viewing the dark-universe  as a skeleton of our universe based only on the Vacuum. Namely at $t=$-$\infty$, $a(t)=\infty$ and starts to decrease until it reaches $a_{\rm min}$ at time $t_{\rm min}$, then begins to increase back to $\infty$ as $t\to+\infty$. At no time in this evolution is $a=0$, again avoiding geometric curvature singularities. There is no singular Big Bang at any time, rather at time $t_{\rm min}$ the universe is hottest since the redshift $z$ is at its maximum. The quantity $t_{0}-t_{\rm min}\approx 13.7$ billion years, does not represent the age of the universe, but rather the time elapsed since the scale factor was at its minimum. The age of this model universe is actually infinite, which we find appealing, since the universe does not originate from an incomprehensible singularity at $a=0$. As described above for the dark-universe, the time evolution of the universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing. All the energy of the hot universe at time $t_{\rm min}$ was accumulated during the downswing, thus such a universe was not created out of nothing, since the Vacuum is not nothing; it is something with a ground state energy density that is a source of energy, if not the original source of all energy in the Universe.
@@
The details of this time evolution of $a(t)$ are very sensitive to the value of $\widehat{b}$ due to the exponentials in (66). For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred above from the Hubble tension. First, one sees from (72) that for small $\widehat{b}$ this amounts to a small corrections to $t_{0}-t_{\rm min}$ which is normally associated to the age of the Universe. If our estimate of $\widehat{b}$ in (66) is approximately correct, then this value for $z_{\rm max}\approx 5\times 10^{21}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$. At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (54) may not be exact at all times, since new degrees of freedom can be excited from the Vacuum and go through various phase transitions that modify (54) in relatively short in time epochs. However if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude. This shows that the Planck scale $\ell_{p}$ plays no role in determining $z_{\rm max}$. However a smaller value for $\widehat{b}$ could easily raise $z_{\rm max}$ up to the Planck scale, where quantum gravity effects could modify our findings.
@@
From the perspective on Gravity presented thus far, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model universe. In the standard cosmology, this earlier era is thought to be described by inflation. Inflation is postulated as a way to get from the singularity at $a=0$ to a hot big bang, however it is not yet universally accepted as being a complete theory. It requires invoking new fields such as the inflaton and fine-tuning their parameters for a “slow-roll”, incorporating re-heating, etc. Very large numbers are naturally involved since the singularity at $a=0$ formally corresponds to $z=\infty$. Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a=0$, then transition to a re-heating scale to bring up the temperature to that of a hot big bang, and this rapid expansion makes the Universe flat and homogeneous. The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is. The duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$ depending on the model, where the upper limit can even be significantly larger than $z_{\rm planck}$. It is noteworthy that at the lower end of this range based on lower reheating temperatures, namely where inflation ends at $z_{\rm inflation}=10^{22}$, is close to our $z_{\rm max}=5\times 10^{21}$. Since this inflationary epoch does not actually exist in our model due to the minimal scale factor $a_{\rm min}$, one should question whether inflation is really necessary. At $t=t_{\rm min}$ our model universe is already expanded since $a(t)>a_{\rm min}$, it is very hot, and is already flat with $k=0$. In Section II we explained why the spatial curvature $k$ is naturally zero since we view Gravity as arising from vacuum energy in flat Minkowski space. The value of $a_{\rm min}=e^{-1/\widehat{b}}$ perhaps plays a role equivalent to hypothetical inflation models since it represents roughly 50 e-foldings since $e^{1/\widehat{b}}\approx e^{50}$ for our estimate of $\widehat{b}\approx 0.02$. On the other hand, one must be careful not to discard the successes of inflationary models, in particular their hallmark prediction of quantum density fluctuations that seed future galaxies and has been observed in the CMB. In our model universe, such quantum fluctuations could arise from the field corresponding to the particle with mass $m_{\rm z}$ itself, which should be interpreted as doing the job of the inflaton. Clearly this requires further study beyond the original scope of this article.
VII Consistency with various types of constraints and experiments
In this section we compare the primary aspects of our model to known observations and experiments of various types. We will base our comparison based on our estimate $\widehat{b}\approx 0.02$, however conclusions are rather sensitive to the value of $\widehat{b}$ and could easily be accommodated with a smaller $\widehat{b}$.
VII.1 $H_{0}(z)$ trend for low $z$ Supernovae
Since (65) is only based on the two data points $z=0$ and $1100$, more compelling support for our proposal would be a fit for a range of redshift $z$ that would confirm the trend proposed above:
$\dfrac{H_{0}(z)}{H_{0}(z=0)}=\sqrt{1-\widehat{b}\log(1+z)}=1-\dfrac{\widehat{b}}{2}\,z+\dfrac{\widehat{b}(2-\widehat{b})}{8}\,z^{2}+{\cal O}(z^{3}).$ (73)
This is possible with supernovae since they exist in a range of redshifts. Remarkably this trend has been very recently observed for a large data sample of supernovae in the range of still relatively low $z$, $0.001<z<2.3$, based on data analyzed at the National Astronomical Observatory of Japan Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 . In the latter work, a fit to the functional form
$\dfrac{H_{0}(z)}{H_{0}(z=0)}\approx\dfrac{1}{(1+z)^{\alpha}}=1-\alpha\,z+\dfrac{\alpha(1+\alpha)}{2}\,z^{2}+{\cal O}(z^{3})$ (74)
was considered, motivated by some versions of $f(R)$ modified gravity and other theoretical models which typically invoke additional fields, for instance Brans-Dicke scalars. (See Montani ; Montani2 ; Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 and references therein.) Although the two formulas (73) and (74) look rather different, to lowest order in small $z$, they agree precisely with $\alpha=\widehat{b}/2$. In Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 values of $\alpha\approx 0.01$ were reported, which agrees with our estimate of $\widehat{b}\approx 0.02$ inferred from the single CMB data point $z=1100$ in (65). This strongly suggests that the trend in $H_{0}(z)$ seen in supernovae extends all the way to $z=1100$, and this provides the strongest support for our model thus far. [^21] We are currently working with Maria Dainotti’s group at NAOJ to perform a detailed statistical analysis in order to try and distinguish between the two formulas (73) and (74). Preliminary results appear to validate our estimate of $\widehat{b}\approx 0.02$ ALMaria . It clearly would be very interesting to extend this range of analysis to supernovae with higher $z$ or to gamma ray bursts.
VII.2 Known bounds on the time-variation of Newton’s constant
If the energy scale varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements. For cosmology, our formula (51) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is time dependent if one ties $\mu$ to the time-dependent scale factor as in (54). Unfortunately, presently there only exists experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$ today, rather than non-null results. Nevertheless, let us check consistency with some known bounds. Based on (55) one has the simple formula
$\dfrac{\dot{G}_{N}}{G_{N}}(t)=H(t)\cdot\widehat{b}$, (75)
where as above $H=\dot{a}/a$. At the present time $t_{0}$,
$\dfrac{\dot{G}_{N}}{G_{N}}=H_{0}\cdot\widehat{b}=6.7\times 10^{-11}\cdot\widehat{b}~/{\rm yr}~~~~({\rm today}),$ (76)
where we have used $H_{0;{\rm CMB}}$. Henceforth, $\dot{G}_{N}/G_{N}$ refers to the present time.
Pulsar timing and gravitational waves.   The strongest experimental constraints come from pulsar timing Taylor based on pulsar observations spanning 30 years Taylor . Over this rather narrow window, one can constrain $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$. These pulsars exist at relatively low redshift $z$. Thus based on equation (76) with $\widehat{b}=0.02$, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars. LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO .
Big Bang Nucleosynthesis    The above results indicate that our model is not yet ruled out up to the CMB scale at $z\approx 1100$. This is still a relatively low energy scale corresponding to a temperature 30003000K. As stated in the Introduction, we assume that (73) extends to higher $z$, and this necessarily involves indirect inferences and consequently will be less conclusive. The next highest energy scale that can provide constraints is Big Bang Nucleosynthesis (BBN) at $z\approx 10^{9}$, corresponding to an energy of about 11  MeV and a temperature of about $10^{9}$K, whose consequences can be inferred from abundances of light nuclei such as Li. Based on the formula (55), at this relatively high $z$, $G_{N}$ is reduced by about 40% if $\widehat{b}=0.02$. The constraints on the time variation of $G_{N}$ for BBN are typically formulated as $\Delta G_{N}/G_{N}$, where $\Delta G_{N}$ is a bound on how much $G_{N}$ can differ at the time of BBN. It’s important that the reported bounds assume slow or no variation after BBN, and $\dot{G}_{N}/G_{N}$ is based solely on a linear fit between the time of BBN and today, and this should be kept in mind in drawing conclusions. Incorporating the full time evolution based on (55) could significantly alter change these bounds. Nevertheless, let us proceed with some simple checks. The earliest article is Krauss concludes that to $1\sigma$ confidence level, $0.85<G_{\rm BBN}/G_{N}<1.21$, which allows a 20% change of $G_{N}$ in either direction in the BBN epoch, which is borderline consistent with our 40% decrease; as we stated a small change in $\widehat{b}$ could potentially accommodate this. To compare with the prediction (76), one should convert this to a time variation over the intervening time, and the same article reports $\dot{G}_{N}/G_{N}<4\times 10^{-13}\,{\rm yr}^{-1}$ which potentially conflicts with $1.4\times 10^{-12}\,/{\rm yr}$ based on $H_{0}$ and $\widehat{b}=0.02$. More recently, tighter bounds were reported by inferring the consequences of BBN on the CMB Alvey . With $2\sigma$ confidence level, this article reports $0.94<G_{\rm BBN}/G_{N}<1.05$, which translated to a time evolution one obtains $\dot{G}_{N}/G_{N}<4.5\times 10^{-12}\,{\rm yr}^{-1}$ Alvey . The latter is consistent with our $1.4\times 10^{-12}\,/{\rm yr}$ estimate above. We conclude that more work is needed here to determine whether our model is consistent with models of BBN, in particular the effect of the complete time evolution which incorporates the time varying $G_{N}$ implicit in (63).
VII.3 Before Big Bang Nucleosynthesis?
At much higher $z$, experimental probes are more severely limited and indirect. Nevertheless some observations are worthwhile to point out. As previously discussed, if the Hubble tension trend extends to this very early universe, then based on our estimate of $\widehat{b}\approx 0.02$, there is a minimal scale factor $a_{\rm min}$ and a corresponding $z_{\rm max}\approx 5\cdot 10^{21}$, which corresponds to a hottest temperature $T_{\rm max}\approx 10^{22}$K, which we associated with a hot big bang. First of all, at least this $T_{\rm max}$ is above the electro-weak scale of about $160$ GeV which is a temperature of about $10^{15}$K, otherwise the electro-weak transition would have never occurred. The temperature $T_{\rm max}$ is roughly the energy scale of some Grand Unified Theories, depending on the model. In the last section we explained how $z_{\rm max}$ is roughly the scale for low-scale inflation models, however our model does not exist for scale factors $0<a(t)<a_{\rm min}$, which is normally considered the inflationary era. This led us to suggest that current models of inflation may not be necessary, since at $z_{\rm max}$ the Universe is already expanded, hot, and flat. Furthermore, before the time when $a(t)=a_{\rm min}$, our solution to $a(t)$ is valid due to (61), namely $a(t)=a(-t+2t_{\rm min})$. A further constraint on our model is that it should not spoil other positive predictions of inflation, in particular how quantum density fluctuations of the hypothetical inflaton field can explain the primordial density fluctuations which are thought to provide the seeds for large scale galaxy formation and probed by the CMB. Quantum fluctuations of the field for lightest particle of mass $m_{\rm z}$, upon which our formula for $\rho_{\rm vac}$} in (2) is based upon, could potentially play such a role, but such considerations require further investigation.
VII.4 Bench-top experiments on the temperature dependence of Newton’s constant
Above we considered implications for cosmology based on the $\Lambda{\rm CDM}$ model and beyond. Apart from the trend recently observed for supernovae discussed above in subsection A, unfortunately we could only compare with bounds rather than definitive non-null experimental signatures. The formula (55) implies that Newton’s constant actually depends on temperature. An analogy can be made with QED where one must take into account the RG flow of the fine structure constant to make predictions at higher energies. For the $\Lambda{\rm CDM}$ model, the energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K. For cosmology, the temperature $T$ as described above is the temperature of the entire Universe, thus it is not clear if such a temperature dependent Newton’s constant is also valid locally in space or time. If it is, and our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (55) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature. If this could be confirmed, it would be truly remarkable that the new fundamental parameter $\widehat{b}$ introduced in this article based on the Hubble tension could be measured in a small bench-top experiment. Ideally one wishes to measure the gravitational force between two masses as a function of their equal temperatures, however we found no such such experiments in the literature. We did find 3 not so well-known experiments wherein the weight of a sample is measured as a function of temperature Shaw ; Dmitriev ; Liangzao , the earliest going as far back as 1923, wherein a positive result is reported rather than a bound. Since only the temperature of the sample is varied, and not that of the Earth, it’s again not clear if our proposal applies directly as stated, however let us proceed.
Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$, i.e. $T=\widetilde{T}+\Delta T$. From (51)
${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)={\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T+\dfrac{\widehat{b}}{2\widetilde{T}^{2}}\,(\Delta T)^{2}+\ldots\right),$ (77)
[…]
VIII Closing Remarks